OpenStudy (anonymous):

16t^2-9=0 solve the equation. round the solutions to the nearest hundredth, if necessary.

OpenStudy (anonymous):

t^2 = 9/16 t = sqrt(9) / sqrt(16) = answer

OpenStudy (anonymous):

\[16t ^{2}-9=0 solve the equation.round the solutions \to the nearest hundredth, if necessary.\]

OpenStudy (anonymous):

t=3/4

OpenStudy (anonymous):

thanx but how you didn't round it